A 10 mg of methyl chloride vapor is added to one liter of nitrogen at 20 C. Pressure is kept constant at 1 atmosphere. What is the final volume and the methyl chloride concentration in mg/cu m and ppmv? ______ Liters; _________mg/m^3 ___________ppmv
1
Expert's answer
2012-04-03T11:26:51-0400
PV = nRT

n (N2) = PV/RT = 101,3 * 1/ 8,31 8(273+20) = 0.0416 mol
n (CH3Cl) = m/Mw = (10 /1000)/50.5 = 0,0002 mol

PV = nRT
V = nRT/P = (0.0416+0.0002)*8,31*(273+20)/101.3 = 1.005 L

w(CH3Cl) = 0.0002/(0.0416+0.0002) = 100% = 0,5%

V = 1,005L * 0.5% /100% = 0.005 L = 0.000005 m3
Density 10/0.000005 = 2000000 mg/m3
ppmv = 5